A right triangle with an area of x^2-4 square units has a leg that measures 2x+4units. Determine the length of the other leg of the triangle.
Answer by nerdybill(7384)

A right triangle with an area of x^2-4 square units has a leg that measures 2x+4units. Determine the length of the other leg of the triangle.
.
Let y = length of the other leg...
then
since area of a triangle is:
(1/2)bh
we then have
x^2-4 = (1/2)(2x+4)y
(x-2)(x+2) = (1/2)2(x+2)y
(x-2)(x+2) = (x+2)y
x-2 = y
.
answer: the other leg is x-2 units